Profit after tax of Starscream Trade Ltd. at the end of the year 2018 is $90,000, some other financial ratios of the year are in the below table:

Return on assets (ROA) =15%

Asset turnover ratio (times)=3

Gross profit margin = 30%

The cost of goods sold for the company in 2018 is:

(a) $1,300,000

(b) $1,620,000

(c) $1,260,000

(d) $1,250,000
Answer #1

Solution:

Return on Assets = Net income / Average Assets

15% = $90000 / Average Assets

Average Assets = $90000 / 15% = $600,000

Now,

Asset Turnover ratio = Sales / Average Assets

3 = Sales / $600,000

Sales = $600,000 * 3 = $1,800,000

Now,

Gross Profit = Sales *Gross Profit margin = $1800000 *30% = $540,000

Cost of goods sold = Sales - Gross Profit = $1800000 - $540000 = $1,260,000

Hence option "C" is correct.